# EgoBlind: Towards Egocentric Visual Assistance for the Blind People

### v1

@@ B.1 Evaluation
To enhance the reliability of the AI evaluation tool (i.e., GPT-4o mini), we sample 30% of the test data to obtain inference results of Gemini 2.0. We invite human volunteers to score the models’ inference results on this subset. Based on the results of human scoring, we refine the prompts to guide the machine’s evaluation results closer to the human scoring outcomes. Ultimately, we calculate the Cohen Kappa coefficient [18] (for inter-rater reliability) between human and GPT scoring results, which is found to be 0.73, indicating a high degree of consistency between the two. Through heatmap analysis (shown in Fig. 19), we observe that samples where human and machine ratings are consistent accounted for 88% of the total test data. For samples where ratings are inconsistent, we find that most of they are due to an information bias – machine cannot see the videos for answer scoring.
@@
B.2 Other Investigations
We conduct targeted experiments to compare the effects with and without specific prompts for the blind. Tab. 6 shows that designing special prompts for blind users can effectively improve model performances, versus naively prompting the models to answer the questions.
Table 6: Effectiveness of our prompt instructions customized for blind people. If without prompting the models that they are answering questions for assisting blind people, remarkable performance drops are observed almost on all question types.
Method Input Tool Use Information Safety Communication Navigation Resource Average
Gemini2.0 Blind-specific 48.3 / 2.9 57.0 / 3.0 51.1 / 2.8 47.7 / 2.7 63.2 / 3.2 65.2 / 3.4 56.6 / 3.0
[…]
InternVL2-8B Blind-specific 61.1 / 3.3 57.2 / 3.0 48.1 / 2.7 59.3 / 3.1 47.4 / 2.5 58.1 / 3.1 56.5 / 3.0
[…]
@@ B.2 Other Investigations
GPT-4o I want you to act as a blind person. I need help identifying blind people. I will provide you with several pictures each time and one question; These questions are all raised by the blind person in the video from his own first-person perspective in the current scene. Your task is to answer the blind person’s question. The answer needs to be based on the content of the picture and the objective characteristics that the blind person cannot see. If the question cannot be answered, you can say I don’t know. Please generate the response with keys ’question’, ’question_id’, ’prediction’, and ’type’. The question is: {question},Question ID: {question_id},Type: {question_type}. The content of the key ’question’ is the question you received. The content of the key ’prediction’ is the answer to the corresponding question you generated. The key ’question_type’ represents which category the question type you received. Your response should look like this example: {’question’:’Is there a wet floor caution sign near the door?’, ’question_id’: ’v_87_1_2’, ’prediction’:’yes.’, ’type’: ’toolrmation reading’}
@@
Gemini I want you to act as a blind person. I need help identifying blind people. I will provide you with a video each time and one question; These questions are all questions raised by the blind person in the video from his own first-person perspective in the current scene. Your task is to answer the blind person’s question which was posed in the last frame of the video. The answer needs to be based on the content of the picture and the objective characteristics that the blind person cannot see. If the question cannot be answered, you can say I don’t know. Do not include Chinese characters in your response. Please generate the response with keys ’question’, ’question_id’, ’prediction’, and ’type’. The question is: {question},Question ID: {question_id},Type: {question_type}. The content of the key ’question’ is the question you received. The content of the key ’prediction’ is the answer to the corresponding question you generated. The key ’question_type’ represents which category the question type you received. Your response should look like this example: {”question”:”Is there a wet floor caution sign near the door?”, ”question_id”: ”v_87_1_2”, ”prediction”:”yes.”, ”type”: ”information reading”}
Table 7: Prompts for models to perform question answering on EgoBlind.
Table 8: Prompts for generating different types of QAs in EgoBlind.
Question Category General Prompts
Social Communication I want you to act as a blind person. I need help identifying blind people.I will provide you with several pictures each time and the pictures have two characteristics:1. pictures have temporal correlation between them and the order of the pictures represents the order of time. 2.The pictures are all taken from your first point of view, which is equivalent to the picture recorded by the camera hanging on your chest. Your task is to imagine what questions you would ask to meet your needs if you were in the situation of the picture provided. You need to focus on asking questions about the types of tools you use which called ‘communication and interaction’.Below I will give a detailed explanation of ‘communication and interaction’. Communication and Interaction:Describe the interaction between blind people and surrounding people/companions, and observe other people‘s status information for multi-person collaborative activities. For example:Did my companion help me carry my luggage onto the conveyor belt? /Has my guide dog entered the elevator safely? /Who is talking to me? Here’s how you can accomplish the task:1.use your imagination and only ask about ‘communication and interaction’questions;2.Please ask questions about each picture;3.The questions you ask need to be real-time and do not contain any off-site information;4.The questions you ask need to be answered through picture content;5.If you cannot answer based on the picture, you can answer ‘I don’t know’;6.Ask as many open-ended questions as possible, not just ’yes-no’ questions;7.Some of the questions generated need to reflect the characteristics of the video, that is, you need to watch several consecutive frames to get the answer.For example:’What is my companion doing? 8.‘communication and interaction’questions should be based on the real life scenarios of blind people, and the questions must be practical (that is, questions that blind people would really ask in the current situation).9.Please don’t ask bad examples like ‘What is my companion saying to me ?’/‘Is there an announcement about the next station or any other important information?’ Please generate the response with ’timestamp’, ’question’, ’answer’ and ’questoin_type’. The content of the key ’timestamp’ the time value corresponding to the current picture,for example:the first image corresponds to the first element in timestamp list, and so on.The timestamp value can be found here:’f’Timestamp list: time.
Information Reading I want you to act as a blind person. I need help identifying blind people.I will provide you with several pictures each time and the pictures have two characteristics:1. pictures have temporal correlation between them and the order of the pictures represents the order of time. 2.The pictures are all taken from your first point of view, which is equivalent to the picture recorded by the camera hanging on your chest. Your task is to imagine what questions you would ask to meet your needs if you were in the situation of the picture provided. You need to focus on asking questions about the types of information reading which called ’information reading’. Below I will give a detailed explanation of ’information reading’. ”Information Reading:Used to obtain some kind of information you want to know. For example:What is the menu about?/What is my mobile phone model?/I just pressed the elevator button. Could you please tell me which floor I just pressed?” Here’s how you can accomplish the task: 1.use your imagination and only ask about ’information reading’ questions; 2.please ask questions about each picture; 3.The questions you ask need to be real-time and do not contain any off-site information; 4.The questions you ask need to be answered through picture content; 5.If you cannot answer based on the picture, you can answer ’I don’t know’; 6.Ask as many open-ended questions as possible, not just ’yes-no’ questions. 7.Some of the questions generated need to reflect the characteristics of the video, that is, you need to watch several consecutive frames to get the answer.For example:’What is the direction of the escalator in front of me?’” 8.’information reading ’ questions should be based on the real life scenarios of blind people, and the questions must be practical (that is, questions that blind people would really ask in the current situation). Please generate the response with keys, ’timestamp’, ’question’,’answer’and ’questoin-type’. The content of the key ’timestamp’ the time value corresponding to the current picture,for example:the first image corresponds to the first element in timestamp list, and so on.”The ’timestamp’ value can be found here:”f”Timestamp list: time.The first image corresponds to the first element in this list, and so on.The content of the key ’question’ is the question you generated.The content of the key ’answer’ is the answer to the corresponding question you generated.The key ’question_type’ represents which category the question type you generated belongs to.You should all generate ’information reading’,your response should look like this example:’timestamp’:’3000’,’question’:’What is this?’,’answer’:’It is a bag.’,’question_type’:’information reading’.Please generate complete data, complete keys, and key values also need to be complete.Please ask questions from first person perspective.Please ask a question for each picture, and the questions asked for different pictures input at the same time need to be different.
Navigation I want you to act as a blind person. I need help identifying blind people.I will provide you with several pictures each time and the pictures have two characteristics:1. pictures have temporal correlation between them and the order of the pictures represents the order of time. 2.The pictures are all taken from your first point of view, which is equivalent to the picture recorded by the camera hanging on your chest. You need to focus on asking questions about the types which called ’local navigation and positioning’.” ”Below I will give a detailed explanation of ’local navigation and positioning’.” ”Local Navigation and Positioning:Find the correct direction/location/navigation information needed by blind people, focusing on simulating indoor navigation and outdoor local navigation functions. For example:where the toilet is? /Am I facing the gate now? /how to get to the nearest stairway from here?” ”Here’s how you can accomplish the task: 1.use your imagination and only ask about ’local navigation and positioning’ questions; 2.please ask questions about each picture; 3.The ’local navigation and positioning’ questions you ask need to be real-time and do not contain any off-site information; 4.The ’local navigation and positioning’ questions you ask need to be answered through picture content; 5.If you cannot answer based on the picture, you can answer ’I don’t know’; 6.Ask as many open-ended ’local navigation and positioning’ questions as possible, not just ’yes-no’ questions. 7.Some of the ’local navigation and positioning’ questions generated need to reflect the characteristics of the video, that is, you need to watch several consecutive frames to get the answer.For example:’What is the direction of the escalator in front of me?’” 8.’local navigation and positioning’ questions should be based on the real life scenarios of blind people, and the questions must be practical (that is, questions that blind people would really ask in the current situation). Please generate the response with keys, ’timestamp’, ’question’,’answer’and ’questoin-type’. The content of the key ’timestamp’ the time value corresponding to the current picture,for example:the first image corresponds to the first element in timestamp list, and so on.”The ’timestamp’ value can be found here:”f”Timestamp list: time.The first image corresponds to the first element in this list, and so on.The content of the key ’question’ is the question you generated.The content of the key ’answer’ is the answer to the corresponding question you generated.The key ’question_type’ represents which category the question type you generated belongs to.You should all generate ’local navigation and positioning’,your response should look like this example:’timestamp’:’3000’,’question’:’Where is the crosswalk?’,’answer’:’It is in front of you.’,’question_type’:’local navigation and positioning’.Please generate complete data, complete keys, and key values also need to be complete.Please ask questions from first person perspective.Please ask a question for each picture, and the questions asked for different pictures input at the same time need to be different.
Safety Warnings I want you to act as a blind person. I need help identifying blind people.I will provide you with several pictures each time and the pictures have two characteristics:1. pictures have temporal correlation between them and the order of the pictures represents the order of time. 2.The pictures are all taken from your first point of view, which is equivalent to the picture recorded by the camera hanging on your chest. Your task is to imagine what questions you would ask to meet your needs if you were in the situation of the picture provided.You need to focus on asking questions about the types of tools you use which called ’safety warnings’. Below I will give a detailed explanation of ’safety warnings’. Safety Warnings:Help blind people determine whether their environment has physical safety hazards/human factor safety risks.For example: Is there any traffic passing by when crossing the road?/What is the direction of the escalator?/Are there any roadblocks during the walk?/Are there any malicious people nearby? Here’s how you can accomplish the task: 1.use your imagination and only ask about ’safety warnings’ questions; 2.please ask questions about each picture; 3.The questions you ask need to be real-time and do not contain any off-site information; 4.The questions you ask need to be answered through picture content; 5.If you cannot answer based on the picture, you can answer ’I don’t know’; 6.Ask as many open-ended questions as possible, not just ’yes-no’ questions. 7.Some of the questions generated need to reflect the characteristics of the video, that is, you need to watch several consecutive frames to get the answer.For example:’What is the direction of the escalator in front of me?’ 8.’safety warnings’ questions should be based on the real life scenarios of blind people, and the questions must be practical (that is, questions that blind people would really ask in the current situation). Please generate the response with keys ’timestamp’, ’question’, ’answer’ and ’questoin_type’.” The content of the key ’timestamp’ the time value corresponding to the current picture,for example:the first image corresponds to the first element in timestamp list, and so on.”The ’timestamp’ value can be found here:”f”Timestamp list: time.The first image corresponds to the first element in this list, and so on.The content of the key ’question’ is the question you generated.The content of the key ’answer’ is the answer to the corresponding question you generated.The key ’question_type’ represents which category the question type you generated belongs to.You should all generate ’safety warnings’your response should look like this example:’timestamp’:’3000’,’question’:’Is it safe to cross the road now?’,’answer’:’yes.’,’question_type’:’safety warnings’.Please generate complete data, complete keys, and key values also need to be complete.Please ask questions from first person perspective.Please ask a question for each picture, and the questions asked for different pictures input at the same time need to be different.
Tool Use I want you to act as a blind person. I need help identifying blind people.I will provide you with several pictures each time and the pictures have two characteristics:1. pictures have temporal correlation between them and the order of the pictures represents the order of time. 2.The pictures are all taken from your first point of view, which is equivalent to the picture recorded by the camera hanging on your chest. Your task is to imagine what questions you would ask to meet your needs if you were in the situation of the picture provided. You need to focus on asking questions about the types of tools you use which called ’tool use’.Below I will give a detailed explanation of ’tool use ’. tool use:Describe how to use a tool or operate a program. For example: How to use microwave oven?/How to use induction cooker and dishwasher? /How to operate the gate? Here’s how you can accomplish the task:1.use your imagination and only ask about ’tool use’ questions;2.please ask questions about each picture;3.The questions you ask need to be real-time and do not contain any off-site information;4.The questions you ask need to be answered through picture content;5.If you cannot answer based on the picture, you can answer ’I don’t know’;6.Ask as many open-ended questions as possible, not just ’yes-no’ questions.7.Some of the questions generated need to reflect the characteristics of the video, that is, you need to watch several consecutive frames to get the answer.For example:’How should I do?’8.’tool use’ questions should be based on the real life scenarios of blind people, and the questions must be practical (that is, questions that blind people would really ask in the current situation). Please generate the response with keys ’timestamp’,’question’,’answer’and’questoin_type’.” The content of the key ’timestamp’ the time value corresponding to the current picture,for example:the first image corresponds to the first element in timestamp list, and so on.”The ’timestamp’ value can be found here:”f”Timestamp list: time.The first image corresponds to the first element in this list, and so on.The content of the key ’question’ is the question you generated.The content of the key ’answer’ is the answer to the corresponding question you generated.The key ’question_type’ represents which category the question type you generated belongs to.You should all generate ’tool use’your response should look like this example:’timestamp’:’3000’,’question’:’How to operate the gate?’,’answer’:’go straight ahead.’,’question_type’:’tool use’.Please generate complete data, complete keys, and key values also need to be complete.Please ask questions from first person perspective.Please ask a question for each picture, and the questions asked for different pictures input at the same time need to be different.
Other Resource I want you to act as a blind person. I need help identifying blind people.I will provide you with several pictures each time and the pictures have two characteristics:1. pictures have temporal correlation between them and the order of the pictures represents the order of time. 2.The pictures are all taken from your first point of view, which is equivalent to the picture recorded by the camera hanging on your chest. Your task is to imagine what questions you would ask to meet your needs if you were in the situation of the picture provided. You need to focus on asking questions about the types of tools you use which called ’resource acquisition and others’. Below I will give a detailed explanation of ’resource acquisition and others’. Resource Acquisition and Others:Looking for nearby services facilities or activities that blind people would like. For example: Is there barrier-free blind walk/manual service/a certain brand store ? ”Here’s how you can accomplish the task: 1.use your imagination and only ask about ’resource acquisition and others’ questions; 2.please ask questions about each picture; 3.The questions you ask need to be real-time and do not contain any off-site information; 4.The questions you ask need to be answered through picture content; 5.If you cannot answer based on the picture, you can answer ’I don’t know’; 6.Ask as many open-ended questions as possible, not just ’yes-no’ questions. 7.Some of the questions generated need to reflect the characteristics of the video, that is, you need to watch several consecutive frames to get the answer.For example:’What is the direction of the escalator in front of me?’” 8.’resource acquisition and others’ questions should be based on the real life scenarios of blind people, and the questions must be practical (that is, questions that blind people would really ask in the current situation). Please generate the response with keys ’timestamp’, ’question’,’answer’and’questoin_type’.” The content of the key ’timestamp’ the time value corresponding to the current picture,for example:the first image corresponds to the first element in timestamp list, and so on.”The ’timestamp’ value can be found here:”f”Timestamp list: time.The first image corresponds to the first element in this list, and so on.The content of the key ’question’ is the question you generated.The content of the key ’answer’ is the answer to the corresponding question you generated.The key ’question_type’ represents which category the question type you generated belongs to.You should all generate ’resource acquisition and others’,your response should look like this example:’timestamp’:’3000’,’question’:’Is there a toilet?’,’answer’:’yes, just go straight ahead.’,’question_type’:’resource acquisition and others’.Please generate complete data, complete keys, and key values also need to be complete.Please ask questions from first person perspective.Please ask a question for each picture, and the questions asked for different pictures input at the same time need to be different.
@@
Table 9: Prompts for GPT-4o mini to evaluate MLLMs on EgoBlind.

### v2

@@ B.1 Evaluation
To enhance the reliability of the AI evaluation tool (i.e., GPT-4o mini), we sample 30% of the test data to obtain inference results of Gemini 2.0. We invite human volunteers to score the models’ inference results on this subset. Based on the results of human scoring, we refine the prompts to guide the machine’s evaluation results closer to the human scoring outcomes. Ultimately, we calculate the Cohen Kappa coefficient [50] (for inter-rater reliability) between human and GPT scoring results, which is found to be 0.73, indicating a high degree of consistency between the two. Through heatmap analysis (shown in Figure 13), we observe that samples where human and machine ratings are consistent accounted for 88% of the total test data. For samples where ratings are inconsistent, we find that most of they are due to an information bias – machine cannot see the videos for answer scoring. Our final evaluation prompts are attached in Table LABEL:tab:evalprompt.
@@
Figure 14: Visualization of failure cases. C: Correctness, S: Score. We only show the key answer words for brevity (same for below).
B.2 Benchmark Analysis
Figure 14 shows other common failure cases of the current Video-LLMs towards egocentric visual assistance for the blind. The example in Figure 14(a) demonstrates that the models are hard to perform effective live QA with streaming visual inputs. They often rely on the wrong frames for responses, even though we have explicitly prompted that the question is posed at the last frame (refer to our prompts at Table 9). The example in Figure 14(b) indicates that the models are poor at scene text recognition, as all model fail to answer the 3rd floor which was pressed by the user. The example also shows that the models are weak in performing temporal context reasoning. Because human can easily know the answer as the user moves her hand upward from button "1" to the 3rd button, even though the button "3" is partly blocked by the hand. Finally, the example in Figure 14(c) suggests that the models are struggling in linking the referential words (e.g.,“this") with the visual contents. This also necessitates spatio-temporal context understanding, e.g., following the hand motion and the object it interacted with.
B.3 Other Investigations
We analyze more details about our investigations on single frame input, blind-specific prompt, and instruction tuning with EgoBlind training set.
B.3.1 Single Frame Input
Figure 15 shows that the overall model performances decline when multi-frame input is replaced with a single frame at the question timestamp. However, the degradations are relatively small (e.g., less than 1%). A breakdown by question type reveals that model accuracy drops significantly for questions related to “Tool Use” and “Navigation,” while it actually improves for “Information Reading” questions which account for nearly half of the dataset. As a result, the overall performance drop is mitigated by these gains. Figure 17(a) shows two cases where models fail to answer using only the question frame but succeed when provided with multiple frames. It reveals that multiple frames are helpful when the question-frame is not in good condition for answering, even if the question itself does not inherently require temporal understanding.
(a) InternVL2.5.
(b) Gemini 1.5 Flash.
(c) GPT-4o.
Figure 15: Uniform video sampling vs. single frame (Q-Frame) input at the question timestamp. The overall QA performance (average accuracy) declines slightly when replacing video with a single question frame. Specifically, for questions of “Navigation”, “Tool Use”, and “Communication”, video-level modeling is key for superior performance, while a single frame can better answer question related to information reading.
B.3.2 Blind-Aware Prompt
Figure 16 demonstrates that a blind-aware prompt are key for better performances. With a normal VQA prompt, the performances of all models decline remarkably for answering almost all types of questions. By analyzing the examples in Figure 17(b), we find that the models are either hard to understand the questions, or tend to give objectively correct answers about the visual contents but are helpfulless for the blind users.
(a) InternVL2.5.
(b) Gemini 1.5 Flash.
(c) GPT-4o.
Figure 16: Blind aware prompting vs. Normal VQA prompting. The model performances decline remarkably in answering almost all types of questions when substituting the blind-aware prompt (Table 9) with a normal vqa prompt (Table 8).
Figure 17: Visualization of predictions.
B.3.3 Instruction Tuning
Implementation Details. The fine-tuning procedures are conducted on two NVIDIA A800 GPUs. Most hyper-parameters are kept at their default configurations as specified in the fine-tuning scripts. For InternVL2.5-8B, optimal performance is achieved by increasing the training epochs to 2 while finetuning only the MLP layer. Regarding LLaVA-OV, we employ LoRA with reduced rank dimensionality (r=16) to optimize computational efficiency. Due to constraints of compute resource, we truncate the maximum sampled frames to 16, and cape the model’s maximum sequence length to 8192, and freeze the vision encoder. Other refinements involve extending the number of epochs to 2, increasing gradient accumulation steps to 4, and adjusting weight decay to 0.05. Finally, for Qwen2.5-VL, we find that the best results are achieved by simply training the model 1 epoch while preserving other parameters unchanged.
Analysis. Figure 18 shows that the overall QA performances of the open-source models can be significantly improved by finetuning with EgoBlind training data. However, we find that the improvements are not consistent across different question types. For “Tool Use” questions, finetuning leads to a performance decline across all 3 models. We speculate that “Tool Use” relies more on procedural and instructional knowledge — areas where large language models (LLMs) excel — but requires less visual understanding. Fine-tuning may overwrite this prior knowledge, causing the models to overfit to the limited training data.
(a) Qwen2.5-VL.
(b) LLaVA-OV.
(c) InternVL-2.5-8B.
Figure 18: Instruction tuning with EgoBlind training data can significantly boost the overall QA performances but is detrimental for answering tool-use questions.
Appendix C Limitations
While we take significant challenge to collect the videos and QAs, the data are limited to the mainland of China so far. Also, the training set is not that big compared with other video QA datasets from sighted people, since there is not much egocentric video data from the blind online. However, we have shown the effectiveness and importance of the training data for better performances. Also, we are continuously collecting related data and aim to extend the dataset by cooperating with the blind association, possibly enclosing blind VQA data from all over the world. In addition, the MLLM techniques are evolving rapidly and we may have missed some most recent models for testing. We will setup an evaluation server and maintain a leader-board to trace the technique advancements in egocentric visual assistance for the blind.
Appendix D Societal Impacts
Positive Impacts: EgoBlind is the first VideoQA dataset that benchmarks vision-language research towards egocentric visual assistance for the blind. The research and outcomes can support developing assistive technology for blind and visually impaired individuals. EgoBlind can advance intelligent systems capable of understanding and describing real-world environments from a first-person perspective, thereby enhancing autonomy, safety, and quality of life for blind users. By enabling more natural and interactive communication between the blind users and the assistive system, VideoQA technologies can help bridge accessibility gaps in daily activities such as navigation, information reading, tool use, social interaction and other resource acquisition, thus holding immerse value towards enlightening the blind individual’s life.
Negative Impacts: Despite its significance. Privacy concerns may arise as egocentric recordings inherently capture sensitive and personal information about both the blind user and bystanders. Additionally, biases in training data or system limitations may lead to inaccurate or misleading responses, potentially compromising user trust or safety. Addressing these challenges is essential to ensure the ethical and inclusive deployment of MLLMs in practical visual assistance for the blind.
@@ Appendix D Societal Impacts
Table 8: Normal prompts for models to perform question answering on EgoBlind.
Model General Prompts
InternVL2.5-26B I will provide you with a video each time and one question; Your task is to answer the question which was posed by the people in the last frame of the video. The answer needs to be based on the content of the picture and the objective characteristics. If the question cannot be answered, you can say I don’t know. Do not include Chinese characters in your response. The question is: row[’question’]
GPT-4o I will provide you with several pictures each time and one question; Your task is to answer the question. The answer needs to be based on the content of the picture and the objective characteristics. If the question cannot be answered, you can say I don’t know. Please generate the response with keys ’question’, ’question_id’, ’prediction’, and ’type’. The question is: {question}, Question ID: {question_id}, Type: {question_type}. The content of the key ’question’ is the question you received. The content of the key ’prediction’ is the answer to the corresponding question you generated. The key ’question_type’ represents which category the question type you received. Your response should look like this example: {’question’:’Is there a wet floor caution sign near the door?’, ’question_id’: ’v_87_1_2’, ’prediction’:’yes.’, ’type’: ’information reading’}
Gemini 1.5 Flash I will provide you with a video each time and one question; Your task is to answer the question which was posed by the people in the last frame of the video. The answer needs to be based on the content of the picture and the objective characteristics. If the question cannot be answered, you can say I don’t know. Do not include Chinese characters in your response. Please generate the response with keys ’question’, ’question_id’, ’prediction’, and ’type’. The question is: {question},Question ID: {question_id},Type: {question_type}. The content of the key ’question’ is the question you received. The content of the key ’prediction’ is the answer to the corresponding question you generated. The key ’question_type’ represents which category the question type you received. Your response should look like this example: {{"question":"Is there a wet floor caution sign near the door?", "question_id": "v_87_1_2", "prediction":"yes.", "type": "information reading"}}
Table 9: Blind-aware prompts for models to perform question answering on EgoBlind.
@@
GPT-4o I want you to act as a visual assistant for the blind. I will provide you with several pictures each time and one question; These questions are all raised by the blind person in the video from his own first-person perspective in the current scene. Your task is to answer the blind person’s question. The answer needs to be based on the content of the picture and the objective characteristics that the blind person cannot see. If the question cannot be answered, you can say I don’t know. Please generate the response with keys ’question’, ’question_id’, ’prediction’, and ’type’. The question is: {question},Question ID: {question_id},Type: {question_type}. The content of the key ’question’ is the question you received. The content of the key ’prediction’ is the answer to the corresponding question you generated. The key ’question_type’ represents which category the question type you received. Your response should look like this example: {’question’:’Is there a wet floor caution sign near the door?’, ’question_id’: ’v_87_1_2’, ’prediction’:’yes.’, ’type’: ’information reading’}
@@
Gemini I want you to act as a visual assistant for the blind. I will provide you with a video each time and one question; These questions are all questions raised by the blind person in the video from his own first-person perspective in the current scene. Your task is to answer the blind person’s question which was posed in the last frame of the video. The answer needs to be based on the content of the picture and the objective characteristics that the blind person cannot see. If the question cannot be answered, you can say I don’t know. Do not include Chinese characters in your response. Please generate the response with keys ’question’, ’question_id’, ’prediction’, and ’type’. The question is: {question},Question ID: {question_id},Type: {question_type}. The content of the key ’question’ is the question you received. The content of the key ’prediction’ is the answer to the corresponding question you generated. The key ’question_type’ represents which category the question type you received. Your response should look like this example: {"question":"Is there a wet floor caution sign near the door?", "question_id": "v_87_1_2", "prediction":"yes.", "type": "information reading"}
Table 10: Prompts for GPT-4o mini to evaluate MLLMs on EgoBlind.
@@
Table 11: Example prompts for GPT-4o to generate QAs.
Social Communication
I want you to act as a blind person. I will provide you with several pictures each time and the pictures have two characteristics: 1. pictures have temporal correlation between them and the order of the pictures represents the order of time. 2.The pictures are all taken from your first point of view, which is equivalent to the picture recorded by the camera hanging on your chest. Your task is to imagine what questions you would ask to meet your needs if you were in the situation of the picture provided. You need to focus on asking questions about the types of tools you use which called ‘communication and interaction’. Below I will give a detailed explanation of ‘communication and interaction’. Communication and Interaction: Describe the interaction between blind people and surrounding people/companions, and observe other people’s status information for multi-person collaborative activities. For example: Did my companion help me carry my luggage onto the conveyor belt? /Has my guide dog entered the elevator safely? /Who is talking to me? Here’s how you can accomplish the task: 1. use your imagination and only ask about ‘communication and interaction’ questions; 2.Please ask questions about each picture; 3. The questions you ask need to be real-time and do not contain any off-site information; 4.The questions you ask need to be answered through picture content; 5.If you cannot answer based on the picture, you can answer ‘I don’t know’; 6. Ask as many open-ended questions as possible, not just ’yes-no’ questions; 7. Some of the questions generated need to reflect the characteristics of the video, that is, you need to watch several consecutive frames to get the answer. For example:’What is my companion doing? 8. ‘communication and interaction’ questions should be based on the real life scenarios of blind people, and the questions must be practical (that is, questions that blind people would really ask in the current situation). 9. Please don’t ask bad examples like ‘What is my companion saying to me ?’/‘Is there an announcement about the next station or any other important information?’ Please generate the response with ’timestamp’, ’question’, ’answer’ and ’questoin_type’. The content of the key ’timestamp’ the time value corresponding to the current picture, for example: the first image corresponds to the first element in timestamp list, and so on. The timestamp value can be found here: ’f’ Timestamp list: time.